### v1

# A Scene Representation for Online Spatial Sonification

Abstract
Robotic perception is emerging as a crucial technology for navigation aids, particularly benefiting individuals with visual impairments through sonification. This paper presents a novel mapping framework that accurately represents spatial geometry for sonification, transforming physical spaces into auditory experiences. By leveraging depth sensors, we convert incrementally built 3D scenes into a compact 360-degree representation based on angular and distance information, aligning with human auditory perception. Our proposed mapping framework utilises a sensor-centric structure, maintaining 2D circular or 3D cylindrical representations, and employs the VDB-GPDF for efficient online mapping. We introduce two sonification modes—circular ranging and circular ranging of objects—along with real-time user control over auditory filters. Incorporating binaural room impulse responses, our framework provides perceptually robust auditory feedback. Quantitative and qualitative evaluations demonstrate superior performance in accuracy, coverage, and timing compared to existing approaches, with effective handling of dynamic objects. The accompanying video showcases the practical application of spatial sonification in room-like environments.
Keywords: Spatial sonification, Scene Representation, Gaussian Process Distance Field.
I Introduction
Robotic perception has the potential to see the environment for human beings. Over the last two decades, robotic sensors such as cameras or LiDARs have served as eyes for robots and have become popular for gathering information from unknown environments. For people living with vision disability, auditory perception is required and could be fulfilled by transforming the information accumulated by the robotic perception via sonification. One of the key steps in representing an environment for sonification is to convert spatial data into an organised format that can be understood through sound. This process facilitates the transformation of physical spaces into auditory experiences, allowing users to understand and navigate through sound. The primary goal of this work is to build a robust framework that can precisely represent an environment’s spatial properties such that it can serve as a basis for sonification.
The ability to incrementally build and maintain a compact representation through sensors will enable real-time auditory feedback to support spatial awareness and environment interaction, especially for visually impaired individuals. As we want our representation to cater towards sonification for humans, we take motivation from how human beings perceive an environment through sound. Thus, we transform a 3D scene captured with depth data into a 360-degree singular representation, based on the angular and distance information. This approach naturally aligns with human auditory perception while simplifying information by focusing on the most relevant spatial points. To achieve this, we propose a mapping framework, based on a sensor-centric organised structure that maintains 2D circular and/or 3D cylindrical rasterised representations (see Fig. 1). These structures capture the geometry of the 3D environment and offer flexibility in how spatial points are sonified by leveraging VDB-GPDF [1], an online efficient mapping framework based on Gaussian Process (GP) distance and gradient fields and a fast-access VDB data structure (Volumetric, Dynamic grid that shares several characteristics with B+trees [2, 3]), enabling a sonifiable online incrementally-built scene representation for indoor and outdoor scenarios.
Figure 1: Our sensor-centric representation for spatial sonification. a) 2D circular representation and b) 3D cylindrical representation for the Cow and Lady dataset.
Moreover, we demonstrate two modes of operation, a circular ranging and a circular ranging of objects sonification modes, which sonify the distance to objects for various directions around the circle. We also introduce various filtering operations to extract information from the proposed 2D circular representation. By incorporating binaural room impulse responses (BRIRs), our mapping framework offers perceptually robust auditory feedback and enhances cognitive mapping. This combination allows for customised sonification of both near-field and far-field objects.
We evaluate quantitatively the performance of the proposed mapping framework in terms of accuracy, coverage and timing, demonstrating the ability to convey the required spatial information better than naively using the current sensor information. Moreover, we show qualitative performance in the presence of dynamic objects and benchmarked with the information provided by distance field mapping. Finally, the submitted video presents the spatial sonification for room-like environments.
II Literature Review
Several studies have explored sonification from vision-based representations. While earlier works focused mostly on 2D vision-to-audio transformations [4, 5], more recent approaches utilise 3D vision-based sensors [6, 7]. Although 3D representations generally offer improved accuracy for localisation tasks, 2D systems can achieve comparable performance for navigation with sonification [8]. Some works have directly used 3D raw data for environment representation [9, 10]. In [9], the authors presented a real-time auditory feedback system that assists blind users in obstacle navigation using a depth camera. The system processes depth data to detect obstacles and sonifies this information. However, using depth data directly can be complex for turning 3D information into clear auditory cues. To tackle this, we propose a 360-degree circular/cylindrical rasterisation method that projects the 3D scenes onto a minimalist geometric structure.
More recently, there has been a growing trend towards leveraging machine learning techniques, such as generative models, to enhance environment representation for spatial sonification. These methods allow for more adaptive and context-aware sonification systems, improving the overall user experience. In [11], the authors introduced “SEE-2-SOUND,” a zero-shot framework. They generated spatial audio from visual inputs such as images, GIFs, and videos without pre-training. A similar task was performed in [12], with additional use of a large language model to generate context-aware synchronised audio-visual content based on the event. While these systems offer adaptability, they often require substantial computational resources. This introduces substantial latency in the sonification process. In contrast, our method provides a lightweight, direct mapping of spatial data to auditory cues. This ensures real-time responsiveness while preserving essential spatial details and the ability to be deployed on modest embedded devices.
There are different techniques to sonify a virtual environment, Gao et al. showed quantitative evidence of how different sonification strategies can influence user perception and behaviour [13]. They compared four elevation sonification mapping strategies and found that binary relative mapping performs better than the others, and in combination with azimuth sonification, it gives clear directional information generating more accurate and efficient task completion. Keeping this in mind, our proposed representation technique is designed to effectively capture both angular and distance information. Thus, ensuring optimal translation of spatial data into auditory cues for the sonification process.
There have been many works on sonification that are dedicated towards visually impaired persons[14, 15]. The type of sound used for sonification also varies depending on the task and the type of information it carries. While discrete sounds are good for simpler information and objects that are near the user, continuous sounds are more suited to deliver complex elements in the scene [16]. Schwartz et al. proposed a mobile application that can reconstruct and sonify a scene to guide users with visual impairment in real-time [15]. The work in [17] built a 3D sonification framework to sonify images from video sequences to send the user a predefined sonic sensorial description through bone-conduction headphones. They combined YOLOV3 [18] for object detection with a probabilistic occupancy map from OctoMap [19] to detect environmental anomalies. These works primarily focus on sonifying immediate surroundings or predefined objects with limited adaptability. They often lack dynamic user control over the sonification process and do not offer full 360-degree spatial coverage, focusing mostly on near-field objects. Our mapping framework, in contrast, is flexible, robust and offers greater adaptability as it conveys spatial information for near-field and far-field.
III Preliminaries
III-A VDB-GPDF Mapping
Our representation is based on the VDB-GPDF mapping framework as proposed in [1]. It couples the VDB grid structure with the Gaussian Process Distance Field [20, 21, 22] to gradually construct and maintain a large-scale, dense distance field map of the environment. The depth sensor data (from depth cameras or LiDARs) at the world coordinate frame is voxelised to create a local VDB structure with 3D points. In every VDB leaf node, the local voxel centres act as training points for GPs. The collection of GPs across all leaf nodes forms the temporary latent Local GP Signed Distance Field (L-GPDF).
For testing, a set of query points is generated by ray-casting from the sensor origin through the voxels in the current reference frame. Each test point is used to query the L-GPDF for the distance field and variance estimates. The queried values from L-GPDF are then fused into a global VDB grid map by updating the voxel distances using the weighted sum method. Following the fusion of the distances and surface properties, the marching cubes algorithm is used to reconstruct a dense surface by identifying the active voxels in the global VDB. The Global Gaussian Process Distance Field (G-GPDF) is made up of the zero-crossing points from active leaf nodes. At this stage, queries for the G-GPDF are computed by averaging the inferred distances from neighboring GP nodes. For further details refer to [1].
IV Methodology
We present an incrementally built scene representation that aims to facilitate spatial sonification of an unknown environment. The proposed representation is based on VDB-GPDF and expanded to cater sonification requirements for human sensory augmentation. As mentioned above, the 3D surface is extracted by applying the Marching Cubes algorithm over the G-GPDF model. propose to simplify this 3D mesh into two distinct rasterisations. First, we squash the 3D surface mesh points into their respective 2D coordinates, and then, these are used to create a 2D circular structure. We also do a simplified 3D cylindrical rasterisation while retaining the full 3D spatial structure. While the 2D circular rasterisation simplifies the 3D environment by projecting the mesh points onto a single 2D circle, the 3D cylindrical rasterisation represents both the horizontal and vertical points on the surface, capturing not only the azimuthal angles but also the height (elevation) of each point. To show the potential of the representation to sonify a 3D environment we develop various strategies for spatial audio feedback, facilitating intuitive sonification with auditory cues.
IV-A 2D Circular Representation
Spatial audio feedback requires accurate but simple information, for this purpose, we propose to represent a dense 3D map of the environment as a 2D circular sensor-centric representation. To ensure our representation aligns with human navigation requirements, we segment ground and non-ground points. Any point that is above the ground and in the height range of a human and can potentially be an obstacle is classified as a non-ground point. Then, we project all these 3D non-ground point cloud or mesh onto a 2D plane. On this plane, every 3D point is reduced to its corresponding 2D coordinates, maintaining spatial relationships within the scene.
A circular grid centred at the location of the depth sensor is created by rasterising these 2D points. Rasterisation is faster and more efficient than raycasting, making it ideal for real-time processes and handling complex scenes. We do this by calculating the radial distance and azimuthal angle for every point in relation to the sensor. The azimuth angle is derived by calculating the cross-product between the sensor’s directional vector and the vector extending from the sensor to the point of interest.
After that, this angle is normalised to fit inside a 360° circular representation. Each degree denotes a distinct angle surrounding the position of the sensor. The distance between the point and the sensor is used to calculate the radial distance. This enables us to rasterise each point, according to its distance from the sensor, onto a circular plane.
This process of rasterisation yields a complete 360° depiction of the scene, with every point corresponding to a distinct angle and distance from the sensor. This ensures capturing the closest visible points surrounding the sensor’s position. Sonifying the three-dimensional scene is especially advantageous with this minimalist circular representation. As we explain in the following section, each point in the circle can be associated with a specific sound, allowing for intuitive sound cues that represent the 3D environment. We ensure that the most pertinent spatial information is preserved by keeping only the closest points at each angle, which improves the audio feedback’s clarity.
IV-B 3D Cylindrical Representation
We take our representation further onto generating a 3D cylindrical structure instead of only a circle. This approach is guided by the intuition that the human is positioned within the cylinder. The reconstructed 3D mesh is rasterised onto a cylindrical surface to create the 3D cylindrical representation. This representation provides a structured, sensor-centric view of the surroundings. The cylindrical grid is discretised into vertical (elevation) and azimuthal intervals. The elevation spans heights between 0.1 and 2 meters above the sensor, while the azimuth is divided into 360-degree intervals.
Each 3D point in the scene is mapped to a specific azimuth-elevation pair, and only the closest point (in terms of radial distance) for each pair is retained. This process effectively reprojects the 3D points onto a simplified 2D representation of the scene. The reprojection occurs by collapsing the radial dimension, projecting all points onto a central 2D plane or line located at the center of the cylinder. The resulting 2D map provides a compact representation of the scene’s geometry, where each point in the azimuth-elevation grid corresponds to the nearest surface point along the cylinder’s radial direction.
IV-C Sonification
We demonstrate the ability of the system’s mapping framework to provide contextual information for cognitive mapping and acquiring spatial information. An auditory sensory augmentation framework is applied in which the information from the spatial representation is rendered or sonified as binaural auditory signals. A foundational aspect of the sonification process is to leverage the natural auditory perspective provided by human spatial hearing. In other words, humans naturally perceive the location of sounds around them and we use the mapping framework to take advantage of spatial hearing perception.
Figure 2: The sound environment used for the BRIR recordings is shown. The HATS manikin is at the lower right.
In this work, augmented or virtual spatial audio is enabled via the use of binaural room impulse responses (BRIRs) which are filters that mathematically specify the transformation of sound from a particular source location in a room to the ears of the participant. There is a separate BRIR filter for each ear and each source location. As well, BRIR filters are unique for each individual in a perceptually relevant way because of morphological differences in ear shape. The data accumulated from the mapping framework provides a representation of the spatial geometry of the room or environment. This information is sufficient to compute BRIRs using room simulation techniques. However, in this work, we use custom BRIRs that were prerecorded in a representative laboratory environment that includes a polypropylene carpet, a motion capture metallic frame, wooden beams with speakers mounted on them and other furniture (see Fig. 2). Significantly, we restrict the source locations to near-field and far-field locations on the audiovisual horizon because source locations on the audiovisual horizon are perceptually more robust in augmented reality applications using generic, i.e., non-individualized, BRIRs.
The BRIR measurements were recorded using the Brüel & Kjær type 4128C Head and Torso Simulator (HATS) with two in-ear microphones together with a Brüel & Kjær type 9640 turntable. We arranged 10 small loudspeakers in a row at a height consistent with the audiovisual horizon of the HATS model. The 10 loudspeakers were positioned directly in front of the HATS model at zero degree of azimuth at distances varying from 0.4 m to 4 m in steps of 0.4 m. Using the turntable, the HATS model was rotated and BRIRs were recorded every $3^{\circ}$ degrees of azimuth from $-177^{\circ}$ to $180^{\circ}$. We also measured a sensor-to-ear transfer function for the Razer Anzu smart glasses and a headphone-to-ear transfer function for the Sennheiser HD 600 open headphones using the HATS model. Appropriate inverse filters were computed for both the smart glasses and headphones for playback of simulated binaural spatial audio signals. The BRIR files were saved as SOFA (Spatially Oriented Format for Acoustics) files [23].
In order to demonstrate the suitability and ease-of-use of the mapping framework for spatial cognitive mapping, we focus on the 2D circular representation. We demonstrate two sonification modes of operation and various filtering operations. One sonification mode, referred to as circular ranging, sonifies the distance to objects for various directions around the circle. It is ‘as if’ the user reaches out horizontally with an imaginary and magical length-changing cane and swings it in a circle counter clockwise. The cane automatically taps the closest object in each direction as it swings about. Interestingly, the range data from the 2D circular representation is discontinuous. There are jumps at the boundary between different objects because objects are separated and their distances vary randomly. We can use this observation advantageously to create an object-based sonification mode. In other words, we can segment the 2D circular range data at the discontinuities forming ‘pseudo objects’ with an angular width defined by the discontinuities. Each object is then sonified only once in the direction and distance of its nearest point as the imaginary cane swings in a circle. We refer to this sonification mode as circular ranging of objects.
For both sonification modes, various filtering operations can be easily applied, providing ‘sensorisensor’ feedback. In other words, the user controls a fundamental aspect of the sensor system – the filtering operation – and receives sensory feedback in the form of changes in the binaural auditory signal. We now describe several filtering operations:
  • field-of-view sonification filter: during the circular operation, sounds are only generated for specified ranges of azimuthal angle.
  • distance sonification filter: during the circular operation, sounds are only generated for specified ranges of distance.
  • object counting filter: during the “circular ranging of objects”, sounds are only generated for a limited number of objects, controlling object clutter.
The details for the demonstration of the two sonification modes are now described. For circular ranging, the 360-degree of range data are separated into 36 angular sectors spanning $10^{\circ}$. For each angular sector, the angle and range corresponding to the smallest range value are selected. The selected range data for each sector is then converted to binaural spatial audio by applying the appropriate BRIR filter to a ‘tap’ sound chosen from the website freesound.org. Ideally, the ‘tap’ sound would relate to material properties, but we leave this for future work and use a constant ‘tap’ sound in this work. Note that if a particular angular sector has not yet been scanned, a ‘woosh’ sound is played instead of a ‘tap’ sound. The distances encompassing the range data are scaled into the 0.4 m to 4 m range of the BRIRs using a constant scaling chosen appropriately for the given environment. This sonification distance range fits well with the perceptual properties of human distance perception. This is to say, human auditory distance perception is significantly poorer than human auditory direction perception [24] and we generally perceive distance in the near-field ($<$ 2 meters) better than the far-field ($>$ 2 meters). To further enhance the distance perception we apply a slight pitch shift to the tap sound. ‘Tap’ sounds closer than 1.5 m are shifted down in pitch by four semitones and ‘tap’ sounds further than 2.5 m are shifted up in pitch by four semitones. The selection of these parameters are arbitrary and simply chosen for demonstration purposes. For the circular ranging of objects mode, we define a range discontinuity as a change of range, $\Delta$, with $\Delta>0.5$ meters/degree. We also limit the number of objects in a single circular scan to the 5 nearest objects. Again, these parameters are arbitrary and chosen for demonstration purposes.
V Evaluation
We quantitatively evaluate the proposed representation for spatial sonification in terms of a) efficiency, b) representation accuracy, and c) comparison with Euclidean distance field, and d) qualitatively demonstrate the representation performance with a dynamic object. Our framework is implemented in C++ based on ROS1. All experiments were run on 12th Gen Intel® Core™ i5-1245U with 12 cores.
Figure 3: The efficiency performance of our proposed representation for spatial sonification with respect to different voxel resolutions.
The first dataset is the Cow and Lady[25], which includes fibreglass models of a large cow and a lady standing side by side in a room. The dataset consists of RGB-D point clouds collected using a Kinect 1 camera, with sensor trajectory captured using a Vicon motion system. We use all frames of the dataset, which covers major information of the scene over 360 degrees. We use the ground truth map of the Cow and Lady, which allows us to perform a proper quantitative evaluation for representation accuracy. We run the online mapping in a room with a live RealSense RGB-D camera as our second dataset. The room has a human as a dynamic object. The poses are computed from the VINS [26] localisation framework.
Figure 4: Quantitative comparisons of the accuracy in RMSE on the cow and lady dataset with varying voxel resolutions.
V-A Time Evaluation
We first evaluate the time consumption to show our representation can provide online performance for sonification. Fig. 3 shows the statistical computational time for the Cow and Lady dataset. Note that our framework has the free-space carving methods [1] enabled to update the moving objects in the scene. For the sake of simplicity, we compute the time for all the processes, including fusion, 2D circle, and 3D cylinder. It illustrates that our representation is efficient with varying resolutions for online mapping.
V-B Representation Accuracy
For the accuracy evaluation, we use the raw depth images in the same way as in our framework to compute the 2D circle and 3D cylinder. Note that we apply the ground truth point cloud of the Cow and Lady to calculate the ground truth circle and cylinder. We then quantitatively compare our circle and cylinder representations and depth images with different map resolutions against the ground truth. As demonstrated in Fig. 4, at the beginning of the dataset, all RMSEs are high due to noisy measurements when the drone is taking off. Then the RMSE drops due to more information is collecting along the steady exploration. The RMSE of the depth sensor circle and cylinder is noisy due to the lack of fusion over time. Our accuracy outperforms depth representations in different resolutions. Note that occasionally, the depth sensor RMSE has lower values than ours, and this is because the depth sensor its only calculated in the field of view of the camera. Therefore, compared to GT and ours, only a few points in the depth circle and cylinder have information and are being evaluated for RMSE.
Figure 5: Quantitative comparisons of the coverage on the cow and lady dataset with varying voxel resolutions.
V-C Coverage Over Frames
We compare the coverage of our framework against the depth sensor again. Due to the ground truth having missing areas, we use the fully reconstructed mesh as the benchmark to compute the coverage. Demonstrating in Fig. 5, as the mapping voxel resolution varies from 5cm to 15cm, the coverage of the depth representations remains the same. The depth circle covers below 20%, and the cylinder covers around 10%. Our incremental representation accumulates and fuses measurements as the coverage grows over frames. Note that the coverage gets up and down occasionally due to passing through unexplored areas to gather more information. With a high resolution of 5cm, our circle reaches 100% coverage, and our cylinder covers over 90% at the end.
Figure 6: We map the scene with a live camera online to show the ability to deal with dynamic objects.
V-D Dynamic Moving Object Validation
We aim to show that our representation has the ability to deal with dynamic objects, which is very common in the real-world scenario. Our framework takes poses and points as raw input, so we use localisation frameworks such as VINs [26] to provide proper poses. We move a live camera to map the scene online. In Fig. 6, our 2D circle is fully reconstructed in 360 degrees, and we can visual object projection on the circle for each angle clearly. From Fig. LABEL:dynamic_1 to Fig. LABEL:dynamic_4, there was a human standing beside the sofa and later walked away. Our circle is updated with respect to the latest status of the scene.
Figure 7: a) Our circular representation shows the distance and directions along the ray. b)However, EDF shows the distance and directions to the closest surfaces.
V-E Compare to EDF
We aim to show our representation is suitable for sonification rather than the Euclidean distance field (EDF), which is a well-used representation in path planning and navigation. For the EDF , it usually has a distance value and the gradient of the distance given the point’s location in the space. It always measures the distance and direction from the point to the closest surface. As we show in Fig.LABEL:arrow_ours, our circle has bearing angles pointing 360 degrees, and for each bearing, we have the closest distance along the ray (the ray arrow is shown in blue). However, as we see from Fig.LABEL:arrow_esdf, we query the EDF for each point on the circle, and this gives us the distance and gradient to the closest surfaces. The gradients of distances (shown in blue) on the circle slip around as the closest surface could be in any direction. This is not suitable for sonification as the bearing does not vary sequentially.
V-F Sonification Demonstration
In our accompanying video [^1] https://tinyurl.com/RepresentationSonification, we show the performance of our proposed representation for sonification with both modes, circular ranging and circular ranging of objects, in the room-size environments of our datasets.
VI Conclusion
We propose an online, incrementally built representation to provide spatial auditory information to aid individuals with visual impairments. By leveraging a sensor-centric mapping structure based on depth sensors, we effectively convert 3D environments into 360-degree auditory representations. Our approach, which includes the VDB-GPDF, plane projection, rasterisation and binaural room impulse responses, ensures fast accurate, complete and perceptually robust mapping for sonification. We demonstrate its performance in both static and dynamic environments. The successful implementation and evaluation of this framework highlight its potential to significantly improve the quality of life for visually impaired individuals by providing an intuitive and efficient means of navigating their surroundings. In future work we aim to perform user studies targeting to reduce the cognitive load of the users.
UTS
University of Technology Sydney
RI
Robotics Institute
FEIT
Faculty of Engineering and Information Technology
WiEIT
Women in Engineering and IT
1D
One-Dimensional
2D
Two-Dimensional
2.5D
Two-and-a-Half-Dimensional
3D
Three-Dimensional
GP
Gaussian Process
GPIS
Gaussian Process Implicit Surfaces
MAP
Maximum A Posteriori
MLE
Maximum Likelihood Estimation
ICP
Iterative Closest Point
MVE
Multi-View Environment
OG
Occupancy Grid
CHOMP
Covariant Hamiltonian Optimization for Motion Planning
FIESTA
Fast Incremental Euclidean DiSTAnce Fields
GPUs
Graphics Processing Units
KD
K-Dimensional
Log-GPIS
Log-Gaussian Process Implicit Surfaces
LiDAR
Light Detection And Ranging Sensor
SLAM
Simultaneous Localisation and Mapping
TDF
Truncated Distance Field
EDF
Euclidean Distance Field
SDF
Signed Distance Field
TSDF
Truncated Signed Distance Field
ESDF
Euclidean Signed Distance Field
GPIS-SDF
GPIS with signed distance function
RGB
Red-Green-Blue
RGB-D
Red-Green-Blue-Depth
RMSE
Root Mean Squared Error
CI
Conditional Independent
FP
Forward Propagation
BP
Backward Propagation
D-SKI
Structured Kernel Interpolation framework with Derivatives
SKI
Structured Kernel Interpolation method
D-SKI-CI-Fusion
Structured Kernel Interpolation with Derivatives and Conditional Independent Fusion
MVMs
Matrix-Vector Multiplications method
PDE
Partial Differential Equation
Log-GPIS-MOP
Log-Gaussian Process Implicit Surface for Mapping, Odometry and Planning
RANSAC
Random Sample Consensus
Dynamic-GPDF
Dynamic Gaussian Process Distance Fields